OpenStudy (anonymous):

In some country, the percent of adult blogging has changed within the various age ranges. From 2007 to 2009, the function y=-3x+28 can be used to estimate the percent of adults under 30 who blogged, and the function y=2x+5 can be used to estimate the percent of adults over 30 who blogged. For both functions, x is the number of years after 2007. A. if this trend continues, predict the year in which the percent of adults under 30 and the percent of adults over 30 who blog was the same. The percent of adults under 30 and the percent of adults over 30 who blog was the same in the year______?

OpenStudy (anonymous):

B. use these equations to predict for the year 2012 the percent of adults under 30 who blog and the percent of adults over 30 who blog. In 2012, ____% of adults under 30 blog. In 2012, ____% of adults over 30 blog

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

ok, x is the number of years after 2007. so x is 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that it?

OpenStudy (sleepyjess):

oh, they want 2012

OpenStudy (anonymous):

so y=-3(2)+28 y=2(3)+5

OpenStudy (sleepyjess):

so how many years away from 2007 is 2012?

OpenStudy (anonymous):

5 years

OpenStudy (sleepyjess):

Yep, so for part b, substitute in 5

OpenStudy (anonymous):

ok one moment, i need to work this out real quick

OpenStudy (sleepyjess):

ok

OpenStudy (anonymous):

y=-3(5)+28=13 y=2(5)+5=15

OpenStudy (aum):

For part A, equate the two y equations, solve for x first. Then add x to 2007 to find the year when the y's will be the same.

OpenStudy (sleepyjess):

@aum got to part A before me :)

OpenStudy (anonymous):

so 13% of adults under 30 blog and 15% of adults over 30 blog Now I need to find the year in part A

OpenStudy (aum):

Your answer to part B is correct.

OpenStudy (anonymous):

how do i find the percent of adults under 30 and the percent of adults over 30 who blog was the same in the year _______________

OpenStudy (sleepyjess):

-3x + 28 = 2x + 5

OpenStudy (sleepyjess):

solve for x

OpenStudy (anonymous):

13=15 or was I not supposed to replace x

OpenStudy (sleepyjess):

Don't replace x, we need to solve for x

OpenStudy (anonymous):

so combine like terms?

OpenStudy (sleepyjess):

Move all x's to one side, all constants to the other

OpenStudy (anonymous):

33=5x

OpenStudy (sleepyjess):

Almost, good try. 23 = 5x

OpenStudy (anonymous):

ooohhh you subtracted 5, lol i added, my bad

OpenStudy (anonymous):

ok do i divide 23/5?

OpenStudy (anonymous):

i get 4.6, ohhh so 2007+ 4 years= 2011 ???

OpenStudy (anonymous):

?

OpenStudy (sleepyjess):

4.6 years, so yeah, I would put 2011

OpenStudy (anonymous):

YES, I have been stuck on this question for hours, thank you so much!!! ;D

OpenStudy (sleepyjess):

\(\LARGE\cal\color{#ff0074}{No~Problem!}\\\bbox [10pt, #00ffab,border:5pt solid blue ]{\color{#872AF0}{\huge\cal ~\dagger~\heartsuit sleepyjess~\heartsuit~\dagger}}\)